### v1

# Context-aware knowledge graph framework for traffic speed forecasting using graph neural network

Abstract
Human mobility is intricately influenced by urban contexts spatially and temporally, constituting essential domain knowledge in understanding traffic systems. While existing traffic forecasting models primarily rely on raw traffic data and advanced deep learning techniques, incorporating contextual information remains underexplored due to the lack of effective integration frameworks and the complexity of urban contexts. This study proposes a novel context-aware knowledge graph (CKG) framework to enhance traffic speed forecasting by effectively modeling spatial and temporal contexts. Employing a relation-dependent integration strategy, the framework generates context-aware representations from the spatial and temporal units of CKG to capture spatio-temporal dependencies of urban contexts. A CKG-GNN model, combining the CKG, dual-view multi-head self-attention (MHSA), and graph neural network (GNN), is then designed to predict traffic speed utilizing these context-aware representations. Our experiments demonstrate that CKG’s configuration significantly influences embedding performance, with ComplEx and KG2E emerging as optimal for embedding spatial and temporal units, respectively. The CKG-GNN model surpasses benchmark models, achieving an average MAE of 3.46±0.01 and a MAPE of 14.76±0.09% for traffic speed predictions from 10 to 120 minutes. The dual-view MHSA analysis reveals the crucial role of relation-dependent features from the context-based view and the model’s ability to prioritize recent time slots in prediction from the sequence-based view. The CKG framework’s model-agnostic nature suggests its potential applicability in various applications of intelligent transportation systems. Overall, this study underscores the importance of incorporating domain-specific contexts into traffic forecasting and merging context-aware knowledge graphs with neural networks to enhance accuracy.
Index Terms:
Traffic forecasting; Context knowledge graph; Spatial and temporal context; Urban transportation; Graph neural network.
I Introduction
Urban transportation is significantly influenced by surrounding environments due to the supply-demand dynamics of human mobility in the urban space [1, 2]. This environment-related influence manifests across both spatial and temporal contexts, which can benefit the predictive models for traffic speed [3, 4]. Actually, context awareness has emerged as a valuable tool in intelligent transportation systems (ITS) to adapt to various traffic scenarios by fully utilizing multi-source context datasets [5, 6]. Despite its potential on ITS, developing context-aware models to boost traffic forecast and investigate contextual effects remains under exploration.
Short-term traffic state prediction involves forecasting upcoming traffic conditions on road segments for periods within hours [7]. Accurate and efficient predictions are key tasks for successful ITS implementation and support multiple downstream tasks, such as traffic signal optimization. Numerous elegant approaches have been contributed to traffic state prediction research. They can be generally classified into two categories: deductive models based on traffic theory and inductive models based on data analysis [8]. Based on a well-established theoretical background, the models based on traffic theory aim at developing simulation models to mimic traffic behavior. These models are computationally intensive, making it challenging to use them for real-time estimation or short-term forecasting. On the other hand, data-driven models produce predictions by extracting and mapping relationships in traffic data, using techniques ranging from basic time series analysis to deep learning. A thorough examination of these research endeavors can be found in [9, 10]. Nowadays, there is a significant increase in the availability of not only traffic datasets but also context datasets, which provide information about the surrounding environments of traffic systems. This proliferation of context datasets opens new opportunities to further improve traffic forecasting [3].
Despite the potential benefits of using surrounding information in traffic forecasting, the context datasets carrying related knowledge are diverse and complex, deepening the difficulty in integrating them into the prediction task [11]. Generally, urban contexts influencing the traffic system encompass various spatial and temporal factors with a broad scope. From a spatial perspective, the morphological layout of a city impacts how transportation systems are designed and operated [2, 12], which in turn affects human mobility and traffic patterns. Spatial datasets, such as points of interest (POIs), land uses, and transportation networks, are typical examples that describe the spatial dimension of urban contexts [4, 13]. Temporally, the traffic patterns vary significantly throughout time periods, weather conditions, and traffic status [7]. For example, peak hours, typically during morning and evening commutes, require a transportation system capable of handling a high volume of passengers. Conversely, during off-peak hours, the system still needs to operate efficiently but with potentially less demand. Regardless of multiple sources, modalities, and dimensions, these spatio-temporal datasets are vital in shaping the nature of urban transportation, becoming a pivotal role in traffic speed prediction [3]. Therefore, effectively organizing context datasets is the prerequisite to incorporating them into the prediction task.
In an effort to organize and facilitate the assimilation of comprehensive traffic-related information, knowledge graphs (KG) emerge as a valuable tool that has attracted substantial attention in the transportation field [14, 15, 16]. The capacity to model relations and domain information makes knowledge graphs powerful in extracting high-level representations from different context datasets by utilizing embedding techniques [17]. However, applying knowledge graphs to traffic forecasting encounters two substantial challenges, i.e., (1) constructing context-aware knowledge graphs associating traffic-related contexts spatially and temporally, and (2) embedding the constructed knowledge graphs and generating context-aware representations spatially and temporally. These two challenges are further deepened considering the diversity and complexity of context datasets. In addition, deep learning architectures, such as graph neural networks (GNN), have become mainstream in traffic forecasting [11, 7, 4], but integrating context representations derived from knowledge graphs with these deep learning models to enhance prediction accuracy is still under exploration. Meanwhile, integrating knowledge graphs into GNNs introduces additional parameters, potentially increasing the model’s complexity, but the domain-specific insights provided by the knowledge graph can enhance the model’s learning efficiency. This can be beneficial in scenarios with limited traffic data, where the structured information from the knowledge graph helps the model to make more informed predictions, leveraging contextual understanding to compensate for data scarcity.
To address these challenges, we propose a context-aware knowledge graph (CKG) framework to effectively model and embed the spatio-temporal relationships inherent in diverse urban contexts. This framework facilitates the integration of various context datasets into machine-readable formats. Then, we strategically combine the proposed CKG framework with GNN through dual-view multi-head self-attention (MHSA) techniques to forecast traffic speed. This integration ensures that CKG-based context representations can be effectively incorporated into traffic forecasting models, leveraging the strengths of both knowledge graphs and advanced neural network architectures to improve predictive performance. Overall, the contributions of this study are three-fold:
  • Propose a context-aware knowledge graph (CKG) framework to model the spatio-temporal relations of context datasets using domain-specific information in the transportation field;
  • Design a relation-dependent integration strategy to generate context-aware representations for traffic speed forecasting. Also, we investigate the performance of various KG embedding techniques and parameter configurations in modeling spatial and temporal contexts;
  • Introduce a dual-view MHSA method to integrate the proposed CKG framework with GNN for traffic speed prediction, i.e., CKG-GNN, assisting in understanding how contexts affect traffic forecasting from the contextual and sequential views.
II Related work
II-A Knowledge graphs in transportation
Knowledge graphs refer to multi-relational graphs linking various entities through relations, which can structure complex datasets to represent and manage knowledge across domains effectively [18, 19]. These graph-based structures enable the modeling of domain-specific knowledge by organizing datasets into a network of interconnected entities and their relations, thus facilitating comprehensive semantic representations [20]. Hence, knowledge graphs are invaluable in applications ranging from ITS to smart city initiatives, where they contribute significantly to the modeling of urban dynamics and environments [21, 15]. For example, projects like the CARES program have demonstrated the potential of integrating dynamic geospatial knowledge graphs with semantic 3D city databases and agent-based systems for intelligent automation [22].
The application of knowledge graphs in the transportation sector has particularly focused on two promising avenues. The first avenue is to establish urban knowledge graphs of movement datasets to capture the spatio-temporal relationships of human mobility directly [23], such as converting the task of human mobility prediction into a knowledge graph completion problem. In detail, Wang et al. [14] introduced a spatio-temporal urban knowledge graph (STKG) to extract structured knowledge from massive trajectory data and then predicted human’s future movement based on the extracted knowledge from spatio-temporal mobility patterns. Second, we can employ knowledge graphs to build relations of context datasets and integrate their representations into traffic forecasting [6, 16]. A representative example is that Zhu et al. [24] constructed a specially designed knowledge graph to encode external factors (e.g., POIs, weather conditions, and time) and then input them into a graph convolutional network to predict traffic speed. Basically, the incorporation of context-aware knowledge graphs into traffic forecasting remains preliminary, with the representation of complex spatio-temporal dependencies posing a significant challenge [24, 16]. Knowledge graph embedding (KGE) techniques provide a promising solution to the challenge by converting logic expressions of entities and relations within knowledge graphs to machine-readable representations [25].
Essentially, KGE techniques map the entities and relations of a knowledge graph into a multidimensional vector space to obtain their vector representations [17, 26]. This process involves defining a scoring function for each fact to ascertain its plausibility, i.e., a triple of (head entity, relation, tail entity), allowing for the quantitative assessment of the relations and entities within the knowledge graph [19]. Wang et al. [20] categorized KGE approaches into two primary types, i.e., translational distance models (distance-based) and semantic matching models (similarity-based). Translational distance models, including TransE and TransR, evaluate the plausibility of facts by measuring the distance between entities after being translated by the relation vector [18]. An adaptation, KG2E, takes uncertainty into account by modeling entities and relations as vectors from multivariate Gaussian distributions [27]. Semantic matching models, on the other hand, rely on similarity-based scoring functions. They determine the plausibility of facts by matching the latent semantics of entities and relations in their vector space embeddings. Examples of this approach include RESCAL, ComplEx, and NTN, which leverage the interplay of vector space dimensions to encode relational patterns [20]. Additionally, several advanced variants of KGE techniques incorporate extra information into the embeddings, such as entity types, relation paths, textual descriptions, and logical rules [18, 20]. These kinds of variants can enrich the representational capacity of knowledge graphs by considering their spatio-temporal dependencies. In this study, we aim to design a context-aware knowledge graph for traffic speed forecasting considering the spatio-temporal relations of available context datasets. Furthermore, we develop a relation-dependent integration strategy to capture the spatio-temporal path dependencies of urban contexts.
II-B Deep learning in traffic forecasting
Traffic forecasting serves as a fundamental component in enhancing transportation resilience, which plays a crucial role in alleviating congestion and facilitating ITS’s development [10]. The scope of traffic forecasting encompasses various indicators, including traffic flow, speed, density, and demand, each offering unique but interconnected insights into the status of transportation networks [7, 28]. Despite different indicators, the goal of traffic prediction is to forecast traffic states based on historical traffic data and external factors. Many methods have been developed to meet this forecasting objective, ranging from conventional statistical approaches to advanced machine learning and deep learning techniques [4]. Notably, deep learning models have gained significant attention for their exceptional ability to handle high-dimensional datasets and capture intricate spatial-temporal dependencies inherent in traffic data [11].
So far, different variants of deep learning models have emerged to forecast traffic speed, including convolutional neural networks (CNN), recurrent neural networks (RNN), and graph neural networks (GNN). CNN-based methods are pivotal for grid-based prediction tasks by leveraging their proficiency in processing spatial data and analyzing traffic patterns within grid-like structures [10]. However, applying CNN-based models to traffic forecasting is constrained due to the graph structure of transportation networks. Meanwhile, the inconvenience of modeling temporal dependencies in sequences also severely limits its performance in the prediction task [4]. In contrast, RNN-based models, such as long short-term memory (LSTM) units and gated recurrent units (GRU), have been proposed to capture the temporal dependencies inherent in traffic data, a feat less attainable by conventional machine learning techniques [10]. Despite their potential in modeling time-series data, RNN-based models fail to elucidate the spatial dependencies in transportation networks since traffic status in a given road is significantly affected by its neighboring roads [4]. The development of GNN-based models represents the forefront of efforts to address the need for capturing both spatial and temporal dependencies in traffic forecasting [7, 28]. Variants such as the diffusion convolution recurrent neural network (DCRNN) [29], spatial-temporal graph convolutional network (STGCN) [30], and temporal graph convolutional network (TGCN) [31] are some typical examples of utilizing GNNs effectively for traffic forecasting. These models underpin the current study’s approach to predicting traffic status, showcasing the advanced capabilities of GNNs in this domain.
Furthermore, urban contexts are also important for traffic forecasting by influencing and shaping human mobility patterns [32, 13]. Tedjopurnomo et al. [11] have underscored the critical role of context information in enhancing traffic forecasting, highlighting the benefits of integrating context-aware insights into predictive models. Nowadays, context-aware modeling is increasingly gaining prominence in the field of traffic forecasting. For example, Zhu et al. [33] proposed an attribute-augmented spatio-temporal graph convolutional network (AST-GCN) to model external factors as dynamic and static attributes for traffic prediction. Zhang et al. [3] specified multimodal contexts from spatial and temporal views, and then proposed a multimodal context-based graph convolutional neural network (MCGCN) model to fuse context datasets into traffic prediction. These examples have verified the superiority of using context information in traffic forecasting compared to existing GNN-based models. Generally, the integration of context-aware strategies into deep learning models represents a pivotal development in traffic forecasting, which lays a robust foundation for this study’s focus on speed prediction by leveraging knowledge graphs.
III Methodology
III-A Problem statement
Assuming $V$, $S$, and $T$ represent raw traffic speed at road segments, spatial context datasets, and temporal context datasets, respectively, the traffic forecast problem is defined as predicting traffic speed in the next $b$ time slots utilizing $V$, $S$, and $T$ in the past $a$ time slots. To solve this problem, this study proposes a CKG framework for traffic forecasting using GNN, i.e., a CKG-GNN, to construct context-aware knowledge graphs and then incorporate the context-aware representations into traffic speed prediction. As shown in Figure 1, the proposed CKG-GNN model consists of three parts: (1) spatio-temporal knowledge graph construction, (2) relation-dependent knowledge graph embedding and integration, and (3) context-aware traffic forecast using dual-view MHSA.
Figure 1: An overview of the CKG-GNN model for traffic speed forecasting.
III-B Spatio-temporal knowledge graph construction
A knowledge graph is a fact-composed graph linking entities through relations. A fact, i.e., an edge in the graph, is expressed as a triple $(h,r,t)$, where $h$, $r$, and $t$ are the head entity, relation, and tail entity, respectively. All observed facts constitute the context-aware knowledge graph, denoted as $\mathbb{K}$. We construct the CKG $\mathbb{K}$ encompassing two units, i.e., a spatial unit $\mathbb{K}_{S}$ and a temporal unit $\mathbb{K}_{T}$, organizing spatial and temporal contexts influencing traffic forecasting, respectively. By using the CKG, we can represent different contexts with the same format to store domain knowledge for the subsequent traffic forecasting task.
III-B1 Spatial unit construction
The spatial unit $\mathbb{K}_{S}$ organizes spatial context datasets by leveraging the interconnected spatial relations among context entities that influence traffic forecasting. In the transportation domain, spatial contexts affecting traffic status and human mobility are diverse. We select three representative spatial contexts according to existing studies, including POIs, road segments, and land uses [32, 24, 3]. These three factors represent spatial contexts from varied spatial dimensions, i.e., spatially-discrete points, spatially-linked lines, and spatially-continuous planes, respectively [3]. Assuming the set of POIs as $P^{s}=\{p_{1},...,p_{i},...,p_{|P^{s}|}\}$, the set of road segments as $R^{s}=\{s_{1},...,s_{i},...,s_{|R^{s}|}\}$, and the set of land uses as $L^{s}=\{l_{1},...,l_{i},...,l_{|L^{s}|}\}$, we establish CKG’s spatial unit $\mathbb{K}_{S}$ using the following three kinds of facts:
  • $\mathbb{K}_{S}(road)$: $(road,adjacentToRoad,road)$. Given $s_{i},s_{j}\allowbreak\in R^{s}$, this fact represents whether road $s_{i}$ is adjacent to road $s_{j}$. Also, we attached an attribute fact to $\mathbb{K}_{S}(road)$ for enriching the semantic information of road entities, namely $(road,hasFFSpeed,speed)$, which provides the free flow speed of the given road. This expression allows us to model the spatial relationship between roads in the CKG, taking roads’ baseline conditions into account.
  • $\mathbb{K}_{S}(poi)$: $(poi,locatedInBuffer[Dist],road)$ and $(poi\allowbreak,hasType,poiType)$. These two facts associate road $s_{i}$ with different types of POIs in $P^{s}$ considering their distance to $s_{i}$. $poiType$ refers to the POI category. $[Dist]$ is the distance to create a buffer for road $s_{i}$ to identify whether a given POI $p_{i}$ is located in the created buffer, as the distances of POIs to a road affect their context effect on human mobility [13]. Here, we construct two buffer types to investigate this distance effect: $[Dist]$ from 10 to 100 meters with an interval of 10 meters and $[Dist]$ from 100 to 500 meters with an interval of 100 meters. Also, we have an attribute fact for $\mathbb{K}_{S}(poi)$, namely $[road,hasPoi[Type]InBuffer[Dist],poiCount]$, meaning the number of POIs with $[Type]$ in $s_{i}$’s buffer with $[Dist]$.
  • $\mathbb{K}_{S}(land)$: $(land,intersectWithBuffer[Dist],road)$ and $(land,hasType,landType)$. The $\mathbb{K}_{S}(land)$ fact is similar to $\mathbb{K}_{S}(poi)$ except for changing the spatial relation from $locatedInBuffer[Dist]$ to $intersect\allowbreak WithBuffer[Dist]$. Instead, the attribute fact is $[road,\allowbreak hasLand[Type]InBuffer[Dist],AreaRatio]$, meaning the intersected area ratio of land $l_{i}$ with $[Type]$ to $s_{i}$’s buffer with $[Dist]$. $landType$ is the land category.
To further utilize the spatial topology information of the transportation network, we introduce a spatial link fact that captures the adjacency order between any two road segments $s_{i}$ and $s_{j}$ as follows:
  • $\mathbb{K}_{S}(link)$: $(road,spatiallyLink[Order],road)$. The fact establishes a connection between $s_{i}$ and $s_{j}$, where $[Order]$ denotes the number of hops between the two roads, with the range capped at 6. For example, an $[Order]$ of 1 implies direct adjacency between $s_{i}$ and $s_{j}$; an $[Order]$ of 2 indicates that the roads are connected via an intermediate road. This fact is designed as it acknowledges that the traffic condition on a given road is influenced not just by its immediate neighbors but also by roads within a broader adjacency context.
III-B2 Temporal unit construction
The temporal unit $\mathbb{K}_{T}$ exemplifies the relations of temporal context datasets in the traffic system, accounting for their temporal changes. In the transportation domain, various time-sensitive factors play a crucial role in traffic forecasting, such as time indicators that mark variations in traffic speed at different times of the day or on different days of the week [10]. To augment $\mathbb{K}_{T}$, we also include traffic jam factors and weather conditions due to their significant influence on traffic dynamics: traffic jam factors are pivotal for tracking the fluctuating nature of vehicle speed and the immediate effect of traffic congestion, and weather information indicates driving behaviors and conditions [11, 7]. For the set of roads $R^{s}$, we define $T^{t}=\{hour_{t},day_{t}\}$ to represent the hour of the day and the day of the week at time $t$. $J^{t}=\{j^{t}_{1},...,j^{t}_{i},...,j^{t}_{|R^{s}|}\}$ refers to the jam factors for each road segment in $R^{s}$ at time $t$, and $W^{t}=\{w^{t}_{1},...,w^{t}_{i},...,w^{t}_{|R^{s}|}\}$ denotes the corresponding weather conditions for these road segments at time $t$. These factors contribute to the construction of the temporal unit $\mathbb{K}_{T}$ through the following facts:
  • $\mathbb{K}_{T}(time)$: $(road,hasHour,hour)$ and $(road,hasDay,\allowbreak day)$. The facts attach time indicators $t$, i.e., the hour of the day and the day of the week, to the given road $s_{i}$. Their attributes reflect human mobility’s difference in various hours ranging [1, 24] and days ranging [1, 7]. A cosine function method is applied to highlight the periodic characteristics of time indicators, with $cosine(2\pi\times hour/24)$ for hours and $cosine(2\pi\times day/7)$ for days.
  • $\mathbb{K}_{T}(jam)$: $(road,hasJam[PastMins],jam)$. Given a road $s_{i}$, this fact captures its average traffic congestion status over a past duration at time $t$. $[PastMins]$ refers to the past minutes, ranging from 10 to 60 minutes with an interval of 10 minutes. The corresponding attribute $jam$ averages jam factors within $[PastMins]$ minutes. Similar to the spatial buffer in $\mathbb{K}_{S}(poi)$ and $\mathbb{K}_{S}(land)$, $[PastMins]$ represents the time buffer for time $t$. This relation helps understand the recent congestion trends on the road to boost the prediction of future traffic conditions.
  • $\mathbb{K}_{T}(weather)$: $(road,hasTprt[PastMins],tprt)$, $(road,hasRain[PastMins],rain)$, and $(road,has\allowbreak Wind\allowbreak[PastMins],wind)$. $\mathbb{K}_{T}(weather)$ comprises three types of relations, each offering an average of the recent weather conditions within $[PastMins]$ minutes at time $t$, including $tprt$ for air temperature, $rain$ for rainfall, and $wind$ for wind speed. Their attribute values, derived from the nearest weather station data to the given road $s_{i}$, are crucial for assessing the impact of weather on traffic dynamics.
In addition, understanding the cyclic nature of traffic dynamics, manifesting on an hourly, daily, or weekly basis, can enhance traffic speed forecasting [11]. This cyclical nature reflects the routine behavior of human mobility, where traffic patterns tend to repeat at similar times on a daily basis (e.g., peak hours in the morning and evening) and weekly (increased traffic on weekends or weekdays). The hourly link is also included due to its direct influence on traffic conditions in the subsequent hour. To include this domain knowledge in $\mathbb{K}_{T}$, we define a temporal link fact that connects road entities with relevant temporal contexts:
  • $\mathbb{K}_{T}(link)$: $(road,temporallyLink[Temp][Link],[Tem\allowbreak p])$. Given a road $s_{i}$, $temporallyLink[Temp][Link]$ is designed to associate it with a specific temporal context $[Temp]$ through a link type $[Link]$. Here, $[Temp]$ could be time indicators derived from $\mathbb{K}_{T}(time)$, jam factors from $\mathbb{K}_{T}(jam)$, or weather conditions from $\mathbb{K}_{T}(weather)$. The $[Link]$ delineates the temporal link on an hourly, daily, or weekly basis, allowing for a thorough representation of time-related influences on traffic forecasting.
Meanwhile, the structure of $\mathbb{K}_{T}$ remains constant over time, while its attributes evolve due to the dynamics of temporal contexts. This setting avoids the need for re-embedding entities and relations within $\mathbb{K}_{T}$ at each time step in the prediction task, thereby improving efficiency while still capturing the essential temporal dynamics.
III-C Relation-dependent knowledge graph embedding and integration
[…]
III-C1 Knowledge graph embedding
KGE techniques are divided into two categories based on their scoring functions: distance-based and similarity-based models [20]. Distance-based models measure the plausibility of facts within $\mathbb{K}$ by calculating the distance between entities following a relation-specific translation. On the other hand, similarity-based models measure the plausibility by matching the latent semantics of entities and relations within their vector spaces. In our study, we delve into both model types, examining their efficacy in CKG embedding and choosing the most suitable methods for embedding $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$.
[…]
$f_{r}(h,t)=-\left\|\textbf{h}+\textbf{r}-\textbf{t}\right\|^{2}_{2}$ (1)
where $\textbf{h},\textbf{r},\textbf{t}\in\mathbb{R}^{e}$. The score is expected to be higher if $(h,r,t)$ holds. However, TransE’s simplistic approach can struggle with complex relation types like 1-to-N, N-to-1, and N-to-N due to its uniform treatment of entity and relation embeddings.
TransR addresses this limitation by introducing relation-specific spaces $\mathbb{R}^{r}$ [18]. Given a fact $(h,r,t)$, TransR projects h and t to the relation-specific space through a projection matrix $\textbf{M}_{r}\in\mathbb{R}^{e\times r}$, i.e., $\hat{\textbf{h}}$ and $\hat{\textbf{t}}$. This projection enables TransR to effectively model diverse relationship types by computing the scoring function as follows:
$\hat{\textbf{h}}=\textbf{M}_{r}\textbf{h},\,\,\,\hat{\textbf{t}}=\textbf{M}_{r}\textbf{t},\,\,\,f_{r}(h,t)=-\left\|\hat{\textbf{h}}+\textbf{r}-\hat{\textbf{t}}\right\|^{2}_{2}$ (2)
KG2E advances the translational concept further by modeling entities and relations as random variables with multi-variate Gaussian distributions, rather than deterministic points in TransE and TransR [27, 18]. This variant enables KG2E to capture the inherent uncertainties in $\mathbb{K}$, with their representations shown in Equation 3.
$\textbf{h}\sim\mathcal{N}(\boldsymbol{\mu}_{h},\boldsymbol{\sigma}_{h}),\,\,\,\textbf{r}\sim\mathcal{N}(\boldsymbol{\mu}_{r},\boldsymbol{\sigma}_{r}),\,\,\,\textbf{t}\sim\mathcal{N}(\boldsymbol{\mu}_{t},\boldsymbol{\sigma}_{t})$ (3)
[…]
In similarity-based models, RESCAL associates each entity with a vector and each relation with a matrix to model the pairwise interactions between latent factors using the bilinear function [20]. The scoring function of RESCAL is defined through the following form, i.e.,
$f_{r}(h,t)=\sum_{i=0}^{e-1}\sum_{j=0}^{e-1}[\textbf{M}_{r}]_{ij}\cdot[\textbf{h}]_{i}\cdot[\textbf{t}]_{j}$ (4)
where $\textbf{h},\textbf{t}\in\mathbb{R}^{e}$ and $\textbf{M}_{r}\in\mathbb{R}^{e\times e}$. However, RESCAL’s reliance on a matrix for each relation can lead to a surge in parameters, thereby elevating computational costs and storage demands.
ComplEx enhances the modeling capability by incorporating complex-valued embeddings, enabling the effective representation of asymmetric relationships [19]. Given a fact $(h,r,t)$, the obtained h, r, and t are in a complex space, denoted as $\mathbb{C}^{e}$, and the scoring function is then defined as follows:
$f_{r}(h,t)=Real\left(\sum_{i=0}^{e-1}[\textbf{r}]_{i}\cdot[\textbf{h}]_{i}\cdot[\bar{\textbf{t}}]_{i}\right)$ (5)
where $Real(\cdot)$ refers to taking the real part of the complex and $\bar{\textbf{t}}$ is the conjugate of t. This scoring function can capture the interaction between entities and relations, making ComplEx adept at handling asymmetric relations.
NTN (neural tensor network) introduces a tensor-based network architecture for embedding [18]. The entities h and t are initially transformed into vector embeddings, i.e., $\textbf{h},\textbf{t}\in\mathbb{R}^{e}$. Then, they are combined through a relation-specific tensor $\dot{\textbf{M}}_{r}\in\mathbb{R}^{e\times e\times r}$ and mapped to a non-linear hidden layer. Finally, a relation-specific output layer computes the score using Equation 6.
$f_{r}(h,t)=\textbf{{r}}^{T}\text{tanh}\left(\textbf{h}^{T}\dot{\textbf{M}}_{r}\textbf{t}+\textbf{M}_{r}^{1}\textbf{h}+\textbf{M}_{r}^{2}\textbf{t}+\textbf{b}_{r}\right)$ (6)
Here, $\textbf{M}_{r}^{1},\textbf{M}_{r}^{2}\in\mathbb{R}^{e\times r}$ are relation-specific matrices and $\textbf{b}_{r}\in\mathbb{R}^{r}$ refers to bias vectors. Despite one of the most expressive models, NTN’s demand for a substantial number of parameters during training poses challenges in terms of efficiency and computational resource requirements [20].
To compare the difference of embedding $\mathbb{K}$ using different methods, we evaluate their performance on link prediction tasks using the PyKEEN library, where the goal is to predict missing relationships in a fact $(h,r,t)$ [34, 26]. The evaluation is distinguished by the side of the prediction, i.e., (1) right-side: predicting $t$ using $h$ and $r$, (2) left-side: predicting $h$ using $r$ and $t$, and (3) both-side: combining both right-side and left-side evaluation. Here, the mean rank (MR) is used to assess the performance of embedding $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$, which averages the ranks assigned to the correct entities within a prediction list [26]. Its value range is $[1,\infty)$, with lower MR values signifying more accurate embeddings. When ranking the prediction list for MR computation, we employ the realistic rank representing the expected value over all permutations respecting the sort order [34].
III-C2 Relation-dependent integration for context representations
After generating embeddings for each entity and relation in $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$ using KGE techniques, we need to strategically deploy these embeddings to foster context-aware embeddings for the traffic forecasting task. Basically, each road in the transportation network represents a discrete unit encapsulating traffic speed data; thus, the simplest way is to incorporate road embeddings derived from $\mathbb{K}$ with speed data directly for the prediction. Nevertheless, this method falls short of leveraging the full potential of $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$, neglecting the graph structure and rich contextual information embedded within the spatial and temporal surroundings. The challenge lies in harnessing the comprehensive context embodied in CKG, which includes the complex spatial and temporal dependencies among various entities and their associated relations.
To solve this challenge, we propose a relation-dependent integration strategy that includes multi-hop relation path integration and attribute augmentation. This strategy is designed to capitalize on the position of road entities in CKG, thereby leveraging the inherent spatial-temporal relationship and context-specific information encoded within its graph structure. A relation path is defined as a succession of linked relations forming a connective sequence between two entities, formally denoted as $r_{1}\rightarrow\ldots\rightarrow r_{l}$ [20]. This multi-hop format enables us to trace and encode the relational trajectories within $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$, capturing the full spectrum of associations surrounding each road segment. Given a relation path $p=r_{1}\rightarrow\ldots\rightarrow r_{l}$ that adjoins a road segment to an entity $e$, and their respective embeddings notated as $\textbf{r}_{1},\ldots,\textbf{r}_{l}$ and e, we define the relation-dependent embedding $\textbf{e}_{path}$ for $e$ in the following mathematical form:
$\textbf{e}_{path}=f_{[\text{dist}]}(\textbf{e}+\sum_{m=1}^{l}\textbf{r}_{m})+f_{[\text{sim}]}(\textbf{e}\cdot\prod_{m=1}^{l}\textbf{r}_{m})$ (7)
Here, $f_{[\text{dist}]}$ and $f_{[\text{sim}]}$ represent indicator functions activated based on the use of distance-based or similarity-based KGE methods, respectively. $f_{[\text{dist}]}(\cdot)=1$ if a distance-based KGE method is utilized, and $f_{[\text{sim}]}(\cdot)=1$ if a similarity-based KGE method is applied. Otherwise, the indicator function defaults to zero. This bifurcation is due to the differing mechanisms by which distance-based and similarity-based models define their respective scoring functions in Section III-C1.
The other component of the proposed relation-dependent integration strategy is attribute augmentation to enhance the semantic richness of each fact. This augmentation employs attribute facts associated with each fact in $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$, such as $hasPoi[Type]InBuffer[Dist]$ for $\mathbb{K}_{S}(poi)$ and $hasLand[Type]InBuffer[Dist]$ for $\mathbb{K}_{S}(land)$, infusing additional attribute information into the embeddings. For simplicity, we denote the attribute value for relations as $x$ and for entities as $y$. Consequently, the attribute-augmented embeddings for relations are $x_{1}\textbf{r}_{1},...,x_{l}\textbf{r}_{l}$, and for entities, it is $y\textbf{e}$. To ensure the consistency and comparability across attributes, we apply max-min normalization to these attribute values separately, resulting in normalized values $x^{\prime}$ and $y^{\prime}$. Integrating these normalized attributes into the relation-dependent embedding yields the enriched context-aware embedding for $e$:
$\textbf{e}^{\prime}_{path}=f_{[\text{dist}]}(y^{\prime}\textbf{e}+\sum_{m=1}^{l}{x_{m}}^{\prime}\textbf{r}_{m})+f_{[\text{sim}]}(y^{\prime}\textbf{e}\cdot\prod_{m=1}^{l}{x_{m}}^{\prime}\textbf{r}_{m})$ (8)
It should be noted when embedding $\mathbb{K}_{S}$, we refine the representation of spatial contexts by directly linking road segments to $poiType$ and $landType$, bypassing the intermediate entities. This simplification, transforming $road\rightarrow poi\rightarrow poiType$ and $road\rightarrow land\rightarrow landType$ into direct $road\rightarrow poiType$ and $road\rightarrow landType$ associations, streamlines the graph structure. It can enhance the semantic integration of POIs and land uses, enabling more effective utilization of context data for the traffic prediction task.
III-D Context-aware traffic forecast using multi-head self-attention
To integrate the context-aware embeddings derived from the CKG into the traffic forecasting model, i.e., CKG-GNN, a preliminary step involves combining these embeddings into a comprehensive representation. In this process, we concatenate the context-aware embeddings in Equation 8 from $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$ according to the following formulation:
$\textbf{e}^{\prime}_{S}=\text{Concat}\left(\textbf{e}_{(road)_{S}},\textbf{e}^{\prime}_{\mathbb{K}_{S}(road;poi;land;link)}\right)$ (9)
$\textbf{e}^{\prime}_{T}=\text{Concat}\left(\textbf{e}_{(road)_{T}},\textbf{e}^{\prime}_{\mathbb{K}_{T}(time;jam;weather;link)}\right)$ (10)
$\textbf{e}^{\prime}_{fuse}=\text{Concat}\left(\textbf{e}^{\prime}_{S},\textbf{e}^{\prime}_{T}\right)$ (11)
Here, $\textbf{e}_{(road)_{S}}$ and $\textbf{e}_{(road)_{T}}$ represent the road-only spatial and temporal features without relation-dependent integration, respectively. $\textbf{e}^{\prime}_{\mathbb{K}_{S}(road;poi;land;link)}$ refers to concatenating the spatial relation-dependent features of roads, POIs, lands, and spatial links. Similarly, $\textbf{e}^{\prime}_{\mathbb{K}_{T}(time;jam;weather;link)}$ is to concatenate the temporal relation-dependent features of times, jam factors, weather conditions, and temporal links. The combined embedding $\textbf{e}^{\prime}_{fuse}$ serves as the context-aware representation for the subsequent fusion process to capture the spatio-temporal dependencies.
Afterward, we utilize the multi-head self-attention (MHSA) mechanism to fuse embeddings from various contexts. The MHSA operates by executing parallel attention operations, emphasizing the focus on different representation segments simultaneously. It computes outputs by correlating queries with a set of key-value pairs, where queries $Q$, keys $K$, and values $V$ are vectors of equal dimensionality from the same source, namely $\textbf{e}^{\prime}_{fuse}$. The core of the MHSA is represented by:
$\text{MultiHead}(Q,K,V)=\text{Concat}[head_{1},\ldots,head_{H}]\boldsymbol{W}_{0}$ (12)
where $\text{where }head_{i}=\text{Attention}(Q\boldsymbol{W}_{i}^{Q},K\boldsymbol{W}_{i}^{K},V\boldsymbol{W}_{i}^{V})$ (13)
In this study, we propose a novel method to apply MHSA from two views, i.e., context-based and sequence-based views, enabling a dual perspective analysis. Initially, the dual-view MHSA discerns the influence of various context features in $\mathbb{K}$, prioritizing those critical for forecasting from the context-based feature view. Subsequently, the sequence-based view employs the output from the context-based view to highlight the historical sequences vital for temporal prediction, emphasizing time slots that are most predictive for future traffic conditions. In addition, forward masking is employed in the sequence-based view to ensure temporal coherence by restricting attention to past and present data. This operation prevents the model from accessing future information, which is essential to avoid lookahead bias. This dual application of MHSA augments the capacity to leverage both contextual features and historical sequences.
Upon deriving the fused context-aware representations from the dual-view MHSA, they are subsequently integrated into a graph convolutional network to facilitate traffic speed forecasting, i.e., CKG-GNN. Specifically, we leverage traffic speed data alongside the context-aware representations from the preceding $a$ time steps to forecast traffic speed for the forthcoming $b$ time steps. In this study, the diffusion convolution recurrent neural network (DCRNN) is employed as the backbone graph convolution model for traffic speed prediction, as detailed by [29]. It is important to note that while DCRNN serves as the backbone of the CKG-GNN in this context, the methodology is designed to be model-agnostic, allowing for the substitution or integration of alternative time-series prediction models. This flexibility underscores the objective to showcase the efficacy of a CKG-based framework in enhancing traffic forecasting accuracy, thereby demonstrating the potential benefits of incorporating context awareness into predictive models in the transportation domain.
IV Experiments and Result
IV-A Datasets and processing
We sourced the traffic speed dataset from HERE technologies, focusing on Singapore’s Core Central Region (CCR), i.e., the downtown area, due to its significant correlation between context information and traffic flow [13]. To investigate the performance of CKG-GNN in different scenarios, we construct three traffic speed datasets, i.e., (1) $V_{raw}$, 06:00-10:00 for three-week workdays from 07/03/2022 to 25/03/2022, emphasizing the workday traffic pattern in morning peak hours; (2) $V_{work}$, 00:00-23:59 on Wednesday, 23/03/2022, a workday dataset; (3) $V_{rest}$, 00:00-23:59 on Saturday, 19/03/2022, a weekend dataset. The traffic speed datasets were collected every 2 minutes from HERE technologies. To enhance data quality and reduce noise, we aggregated the speed datasets into 10-min intervals by averaging. A demo of raw speed data in CCR is shown in Figure 2.
Figure 2: Overview of Singapore’s Core Central Region (CCR) and the used datasets. The speed shown in the figure is a demo dataset collected from HERE technologies.
To establish $\mathbb{K}_{S}$, we collected three spatial context datasets, including road segments, POIs, and land uses shown in Figure 2. (1) Road segments: The transportation network in CCR is built using shape information from HERE technologies, resulting in 1,606 viable road segments after correcting topological inaccuracies. The dataset also provides road segment lengths and free-flow speed information for creating $\mathbb{K}_{S}(road)$. (2) POIs: The POI dataset aggregates data sources from Singapore OneMap, DataMall, and OpenStreetMap, yielding 17 categories with 85,647 points across Singapore [3]. It encompasses diverse classes, such as community, education, commercial, and residential areas, to form $\mathbb{K}_{S}(poi)$. (3) Land uses: This dataset sourced from Singapore’s Urban Redevelopment Authority in 2019, includes 28 reclassified categories such as business, residential, and park areas, offering a comprehensive view of the urban landscape to construct $\mathbb{K}_{S}(land)$.
Regarding the construction of $\mathbb{K}_{T}$, temporal contexts like time indicators, traffic jam factors, and weather conditions are collected due to their influential impact on traffic dynamics. (1) Time: This context indicates the hour of the day and the day of the week to distinguish the temporal changes of traffic dynamics in composing $\mathbb{K}_{T}(time)$. (2) Traffic jam factor: Collected every 2 minutes from HERE Technologies, this dataset provides real-time congestion levels ranging from 0.0 (free flow) to 10.0 (road closure) for each road segment. We aggregated these values into 10-min intervals to align with the temporal resolution of the traffic speed dataset, and then constructed $\mathbb{K}_{T}(jam)$. (3) Weather conditions: We collected real-time data on air temperature, rainfall, and wind speed from Singapore’s weather stations to build $\mathbb{K}_{T}(weather)$, aligning their temporal granularity with the traffic speed dataset by resampling them to 10-min intervals. Weather data was spatially associated with road segments based on proximity to the nearest weather stations. The distribution of weather monitoring stations is illustrated in Figure 2, and the types of weather conditions collected vary by station, with some stations capturing all three conditions while others collect only one or two, depending on the stations’ settings.
IV-B Spatial KG embedding evaluation
Embedding $\mathbb{K}_{S}$ involves estimating MR’s performance with different KGE models and parameters to select the optimal configuration by capturing the complex interactions of spatial factors. In Table I, distance-based methods generally excel in embedding the spatial unit $\mathbb{K}_{S}$, as indicated by their lower both-side MR values compared to similarity-based methods. However, ComplEx, a similarity-based method, defies this trend by achieving the lowest MR (12.46 for Buffer[10-100] and Link[6]), suggesting its effectiveness in capturing spatial relationships in $\mathbb{K}_{S}$. Exclusively under the conditions of Buffer[10-100] and Link[-], TransE attains the best MR of 50.64. The superior performance of ComplEx is attributed to its ability to model asymmetric relationships through complex-valued embeddings in $\mathbb{K}_{S}$ [19]. Refer to Section V-A for the asymmetric phenomenon of the left-side and right-side evaluation. Also, the result aligns with the notion that similarity-based methods like RESCAL and NTN require a substantial amount of facts for training to perform optimally.
TABLE I: KG embedding performance (both-side MR) of the spatial unit $\mathbb{K}_{S}$ with different configurations of buffer [Dist]. There are two spatial links: Link[-] (no link) and Link[6] ([Order]=6).
Model Buffer[10-100] Buffer[100-500] Buffer[10-500]
Link[-] Link[6] Link[-] Link[6] Link[-] Link[6]
TransE 50.64 20.40 72.89 33.25 70.27 39.44
TransR 58.74 18.23 71.73 32.28 66.88 35.60
KG2E 54.09 21.18 76.38 35.11 72.11 37.77
RESCAL 764.17 824.76 705.22 793.25 720.09 785.08
ComplEx 53.66 12.46 66.52 24.86 57.99 27.88
NTN 429.00 51.28 180.69 55.80 177.33 67.01
In addition, Table I also delves into the impact of buffer distances on KG embeddings for $\mathbb{K}_{S}(poi)$ and $\mathbb{K}_{S}(land)$. Different buffer ranges, namely Buffer[10-100] and Buffer[100-500], are examined alongside a combined range of Buffer[10-500]. We find that Buffer[10-100] achieves the best performance for well-performing methods like TransE, TransR, KG2E, and ComplEx, suggesting that a closer proximity range is sufficient for associating POIs and land uses with the road network for effective embedding. Extending the buffer to include more distant POIs and land parcels may introduce noise, potentially confounding the embedding process. Also, the MR evaluation under two spatial link conditions, Link[-] and Link[6], justifies the observations regarding KGE methods and buffer distance efficacy, underscoring the robustness of the embedding strategies across different spatial link scenarios.
TABLE II: Both-side MR evaluation of various spatial links in the KG embedding of $\mathbb{K}_{S}$. Link[-] is no link, and Link[1,…,6] is [Order]=1,…,6.
Model Buffer Link[-] Link[1] Link[2] Link[3] Link[4] Link[5] Link[6] ComplEx [10-100] 53.66 49.38 41.72 33.26 23.21 17.09 12.46 TransE 50.64 48.10 43.70 37.94 30.81 24.44 20.40 ComplEx [100-500] 66.52 63.36 58.60 50.69 41.04 31.90 24.86 TransE 72.89 69.90 65.62 58.34 49.07 40.36 33.25 ComplEx [10-500] 57.99 57.26 53.97 48.52 42.19 34.78 27.88 TransE 70.27 68.58 64.82 60.65 55.01 47.47 39.44
Given that both ComplEx and TransE demonstrate the potential to achieve the lowest MR in some scenarios in Table I, we further explore their performance across diverse spatial link configurations. As shown in Table II, While TransE outperforms in the specific scenario of Link[-] and Link[1] for Buffer[10-100], ComplEx demonstrates superior performance across all other configurations. A closer examination of the impact of different spatial links reveals a consistent trend: an enhancement in spatial links, as indicated by lower both-side MR values, correlates with improved embedding performance. This trend suggests that denser connections within the transportation network contribute positively to embedding $\mathbb{K}_{S}$, a pattern that holds across all buffer distances. Consequently, the optimal CKG configuration for embedding $\mathbb{K}_{S}$ appears to be Link[6] with Buffer[10-100], where ComplEx achieves the lowest both-side MR, specifically 12.46, implying its potential in providing the most suitable embeddings for subsequent traffic forecasting tasks.
IV-C Temporal KG embedding evaluation
Similar to embedding $\mathbb{K}_{S}$, selecting appropriate methods and parameters to embed the temporal unit $\mathbb{K}_{T}$ is also crucial. Table III demonstrates the both-side MR results for embedding $\mathbb{K}_{T}$ using both distance-based and similarity-based methods. Notably, KG2E consistently outperforms other methods across different configurations with the lowest MR values, benefiting from its approach of representing entities and relations through multi-variate Gaussian distributions [27]. This feature makes KG2E incorporate uncertainties into the embedding process, capturing the inherent variability and complexity of relations in $\mathbb{K}_{T}$, especially in dynamic systems like traffic networks. Conversely, similarity-based methods struggle in sparse temporal connections in $\mathbb{K}_{T}$ because these methods rely on a denser graph structure to uncover and interpret the subtle patterns. Regarding the impact of [PastMins], a distinct pattern emerges: larger [PastMins] values correlate with lower MR values across all methods, suggesting that a broader time buffer enhances embedding effectiveness. The cutoff at Past[60] is strategically chosen to align with the hourly link, and extending the time buffer beyond this point could introduce redundant complexity. Therefore, Past[60] is determined to be the optimal choice. In addition, the above observations hold across two kinds of temporal links (Link[-] and Link[HDW]), reinforcing the findings’ consistency and reliability.
TABLE III: KG embedding performance (both-side MR) of the temporal unit $\mathbb{K}_{T}$ with different configurations of [PastMins]. There are two temporal link situations, i.e., [-] referring to no links and [HDW] including hourly, daily, and weekly links.
Model Past[10] Past[20] Past[30] Past[40] Past[50] Past[60] [-] [HDW] [-] [HDW] [-] [HDW] [-] [HDW] [-] [HDW] [-] [HDW] TransE 6.35 3.97 5.69 3.98 4.98 3.73 3.97 3.79 3.98 3.33 3.73 3.30 TransR 18.93 10.69 18.56 12.90 11.00 13.15 10.69 9.08 12.90 8.49 13.15 7.02 KG2E 5.47 3.49 4.13 3.18 3.75 3.18 3.49 2.85 3.18 2.59 3.18 2.36 RESCAL 282.31 406.15 407.05 405.33 407.25 407.00 406.15 405.54 405.33 406.28 407.00 407.20 ComplEx 388.57 15.49 193.17 7.79 50.09 5.53 20.22 4.36 8.09 4.18 6.07 4.51 NTN 352.73 290.40 349.04 242.79 291.22 225.85 290.40 186.63 242.79 149.17 225.85 123.72
To further investigate the most effective temporal link configuration, Table IV evaluates the both-side MR values for different temporal links with all [PastMins] using KG2E. It is noted that Link[HDW] consistently achieves the best performance, with the lowest both-side MR of 2.36 at Past[60]. Interestingly, the MR values for single temporal links—[hour], [day], and [week]—are identical. This uniformity suggests that when embedding $\mathbb{K}_{T}$, the temporal links alone (hourly, daily, or weekly) do not differentiate connections among road segments as they all evolve over time. However, when integrating all temporal links (Link[HDW]), $\mathbb{K}_{T}$ achieves an optimal structure for embedding. In summary, embedding $\mathbb{K}_{T}$ using KG2E is most effective with Link[HDW] and Past[60], offering the most effective CKG embedding for the subsequent traffic forecasting tasks.
TABLE IV: Both-side MR evaluation of various temporal links in the KG embedding of $\mathbb{K}_{T}$ using KG2E.
Model Link Past[10] Past[20] Past[30] Past[40] Past[50] Past[60]
KG2E [-] 5.47 4.13 3.75 3.49 3.18 3.18
KG2E [hour] 4.13 3.75 3.49 3.18 3.18 2.85
KG2E [day] 4.13 3.75 3.49 3.18 3.18 2.85
KG2E [week] 4.13 3.75 3.49 3.18 3.18 2.85
KG2E [HDW] 3.49 3.18 3.18 2.85 2.59 2.36
IV-D Performance of traffic speed prediction
To investigate the performance of CKG-GNN models in traffic forecasting, we utilized speed datasets $V_{raw}$ for experiments. The results, as shown in Table V, illustrate the evaluation metrics of various forecasting models over time horizons of 10, 60, and 120 minutes. The baseline GNN models establish benchmarks for comparison using the LibCity library, including TGCN, STGCN, and DCRNN [35], while the proposed CKG-GNN models integrate CKG with GNN using different context-aware units. Specifically, CKG-$\mathbb{K}_{S}$ employs the spatial unit, CKG-$\mathbb{K}_{T}$ incorporates the temporal unit, and CKG-$\mathbb{K}_{ST}$ combines both spatial and temporal units. Evaluation metrics, including MAE (mean absolute error) and MAPE (mean absolute percentage error), are presented as averages with standard deviations across five runs to present reliability and insights into the models’ consistency. For the experimental setup, we configured the epoch numbers at 500 with a batch size of 16. The optimization was performed using the Adam optimizer, initiated with a learning rate of 0.001. We applied a MultiStepLR scheduler to enhance the convergence behavior, reducing the learning rate by 0.5 at epochs 150, 250, 350, and 450. For the dual-view MHSA, we set the head numbers for the context view as 10 and for the sequence view as 4, taking into account the feature dimensions associated with each view. Furthermore, the dataset was divided into training, validation, and testing sets with ratios of 0.7:0.1:0.2, respectively.
TABLE V: Performance comparison of predicting traffic speed using different models in 10/60/120 minutes and its average values from 10 to 120 minutes using MAE (km/h) and MAPE (%). Each experiment was run five times to obtain its mean and standard deviation.
Model 10 min 60 min 120 min Ave (10-120 min)
MAE MAPE MAE MAPE MAE MAPE MAE MAPE
TGCN 4.82±0.04 20.66±0.30 5.34±0.03 24.10±0.17 5.52±0.06 25.34±0.31 5.32±0.03 23.92±0.15
STGCN 3.29±0.05 13.93±0.12 3.78±0.05 16.19±0.44 3.76±0.08 16.21±0.61 3.72±0.04 15.95±0.37
DCRNN 3.14±0.00 13.15±0.02 3.69±0.01 16.13±0.07 3.73±0.01 16.47±0.13 3.64±0.01 15.86±0.07
CKG-$\mathbb{K}_{S}$ 3.16±0.00 13.12±0.02 3.65±0.03 15.61±0.11 3.64±0.03 15.94±0.16 3.60±0.02 15.46±0.09
CKG-$\mathbb{K}_{T}$ 3.09±0.01 12.95±0.05 3.56±0.01 15.37±0.09 3.56±0.02 15.39±0.10 3.51±0.01 15.10±0.08
CKG-$\mathbb{K}_{ST}$ 3.08±0.01 12.88±0.05 3.52±0.00 15.04±0.15 3.47±0.02 14.86±0.11 3.46±0.01 14.76±0.09
The results in Table V underscore the significant boost in forecasting performance achieved by integrating the CKG with GNN across various time horizons. When examining the average performance from 10 to 120 minutes, DCRNN stands out among the baseline models with an MAE of 3.64±0.01 and a MAPE of 15.86±0.07%. After introducing the spatial unit $\mathbb{K}_{S}$ to GNN, we observe a notable improvement compared to DCRNN, evidenced by reductions in MAE and MAPE. This enhancement is further amplified using the temporal unit $\mathbb{K}_{T}$, suggesting that temporal factors have a more pronounced impact on forecasting accuracy than spatial factors. This improvement is attributed to the direct impact of temporal factors on traffic forecasting due to their alignment with the dynamic nature of traffic flow. In contrast, spatial contexts contribute by delineating the intricate supply-demand relationship of human mobility over urban spaces, a complex aspect challenging to leverage effectively in traffic prediction models [3]. The best results are obtained when both $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$ are integrated, i.e., CKG-$\mathbb{K}_{ST}$ with the MAE and MAPE of 3.46±0.01 and 14.76±0.09%, respectively. This showcases the combined strength of spatial and temporal contexts in boosting traffic predictions.
The individual performance of 10, 60, and 120 minutes reveals that the CKG-enhanced models consistently outperform the baselines, with a minor exception at the 10-min prediction using only $\mathbb{K}_{S}$. Although the 10-min MAE for $\mathbb{K}_{S}$ is slightly worse than DCRNN, the MAPE still shows an improvement. Furthermore, CKG-$\mathbb{K}_{ST}$ outperforms all the other models from 10-min to 120-min prediction, suggesting that the combination of spatio-temporal contexts contributes to the predictive reliability over both short and longer-term horizons. Regarding error progression from 10-min to 120-min forecasts, baseline models like DCRNN exhibit a gradual increase in MAE, but this error accumulation can be mitigated when context-aware knowledge graphs are integrated. Despite an increase in MAE from 10-min to 60-min forecasts across $\mathbb{K}_{S}$, $\mathbb{K}_{T}$, and $\mathbb{K}_{ST}$, their performance from 60 minutes to 120 minutes remains stable or even improves, such as the MAE decrease from 3.52±0.00 to 3.46±0.01 for CKG-$\mathbb{K}_{ST}$. This stability indicates the potent effect of CKG in enhancing longer-term traffic speed predictions.
IV-E Dual-view feature importance
Figure 3(a) demonstrates the heatmap derived from context-view attention weights in Equation 12 to investigate the contribution of context features. For the spatial unit, the prominence of the relation-dependent road feature at $[1]_{a}$ stands out, surpassing the influence of the road-only spatial feature at $[0]_{a}$ and the road-only temporal feature at $[10]_{a}$. This phenomenon suggests that isolated road representations without additional context information are less predictive. Moreover, the weights in Rows $[4-8]_{a}$ are notably elevated, indicating that spatial links—ranging from proximate to extended connections—play a crucial role in capturing the spatial dependency affecting traffic prediction. These links also exhibit a strong interplay with temporal features, particularly time indicators at Col $[11]_{a}$ and weather conditions at Col $[13]_{a}$.
Figure 3: Attention weight heatmaps for context-view and sequence-view features in CKG-$\mathbb{K}_{ST}$. (a) Heatmap for context-view features with $[0-9]_{a}$ for the spatial unit and $[10-16]_{a}$ for the temporal unit. $[0]_{a}$: road-only spatial feature; $[1-3]_{a}$: relation-dependent features for roads, POIs, and land uses, respectively; $[4-9]_{a}$: spatial links from 1 to 6. $[10]_{a}$: road-only temporal feature; $[11-13]_{a}$: relation-dependent features for time indicators, jam factors, and weather conditions, respectively; $[14-16]_{a}$: temporal links for hour, day, and week, respectively. (b) Heatmap for sequence-view features in the last 12 time slots with masks. $[0]_{b}$: the earliest; $[11]_{b}$: the latest.
Regarding the temporal unit, all features within $[11-13]_{a}$ manifest as critical. First, time indicators and weather conditions in Rows $[11,13]_{a}$ demonstrate substantial weights individually and mutually. Their interaction with the relation-dependent road feature at Col $[1]_{a}$ is also of great importance. Second, the jam factor at $[12]_{a}$ stands out for its significant self-weight, aligning with the intuitive understanding that the congestion level is directly predictive of traffic speed with an immediate impact. Furthermore, hourly and weekly patterns in Rows $[14,16]_{a}$ carry high weights, whereas daily patterns in Row $[15]_{a}$ appear less influential. The prominence of hourly patterns can be attributed to their immediate relevance to the prediction timeframe. Meanwhile, weekly patterns demonstrate consistent regularity, capturing the cyclical nature of traffic flow where behaviors on a specific day, like Monday, tend to mirror those from the previous week. In contrast, daily patterns exhibit a weaker correlation, suggesting that the connection between consecutive days, such as Friday to the following Monday, possesses limited predictive value in this context.
On the other hand, Figure 3(b) provides an attention-weight heatmap from the sequence view over the last 12 time slots, where $[0]_{b}$ denotes the earliest and $[11]_{b}$ denotes the latest time slot. The upper-right corner of the heatmap is filled with zeros, a direct result of employing MHSA masks to prevent the model from accessing future information, ensuring that predictions are based solely on past and present data. The heatmap illustrates that cells along and near the diagonal possess higher weights, underscoring the model’s recognition of more immediate past events as more indicative of the current traffic speed. For instance, in Row $[1]_{b}$, the weight at Col $[1]_{b}$ is greater than that at Col $[0]_{b}$, aligning with the self-attention mechanism’s design to assign more importance to proximate temporal information. Moreover, the heatmap presents a gradient of diminishing weights progressing downward from Row $[0]_{b}$ to $[11]_{b}$, illustrating a gradual reduction in the influence of historical time slots on the present prediction. Overall, the heatmap thus not only visualizes the sequential dependencies within the time-series data but also reveals the diminishing impact of distant past information on current traffic speed forecasting.
V Discussion
V-A CKG embedding analysis
In Sections IV-B and IV-C, we have analyzed the both-side evaluation of CKG embeddings for $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$. A further exploration into the left-side (head prediction) and right-side (tail prediction) evaluations can unveil additional insights, as shown in Figure 4. In detail, Figure 4(a, b) employs ComplEx and TransE methods across diverse spatial buffer ranges and spatial links, while Figure 4(c, d) leverages KG2E and TransE methods to explore different time buffers and temporal links.
Figure 4: Embedding performance of both-side (b), left-side (head, h), and right-side (tail, t) link predictions for $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$. (a) ComplEx for $\mathbb{K}_{S}$. (b) TransE for $\mathbb{K}_{S}$. (c) KG2E for $\mathbb{K}_{T}$. (d) TransE for $\mathbb{K}_{T}$.
Regarding $\mathbb{K}_{S}$, we have several findings. (1) Head and tail predictions. An imbalance in MR values for left-side and right-side evaluations is observed for both methods in Figure 4(a, b), with the both-side case presenting intermediate values. This imbalance reveals that predicting the head entity (e.g., $poiType$ and $landType$) given the tail entity and the relation is more straightforward compared to the reverse. We infer that the asymmetry stems from the more distinguishable spatial distribution of POI and land types, making them easier to associate with specific road segments compared to predicting a road based on its surrounding urban functions. (2) Spatial link influences. A decline in MR with increasing link numbers suggests that additional spatial links bolster the $\mathbb{K}_{S}$’s ability to delineate interactions and relationships between roads, POIs, and land uses. To further explore the link influence, we also investigate the embedding performance for Link[9] and Link[12] by adding more spatial links. However, the inclusion of Link[9] and Link[12] does not yield a large performance enhancement beyond Link[6]. For TransE, the performance on the right-side prediction even deteriorates for Link[9] and Link[12]. Hence, Link[6] is recommended for an optimal parameter of $\mathbb{K}_{S}$, balancing performance gains and computational resource usage. (3) Spatial buffer selection. We find that Buffer[10-100] achieves the best MR scores across three-side predictions using both models, corroborating findings from Section IV-B. This particular buffer range encapsulates essential spatial relationships in $\mathbb{K}_{S}$ among entities, striking a balance where the context is rich without being overly broad.
Similarly, we analyze the three-side MR values for $\mathbb{K}_{T}$ using KG2E and TransE methods in Figure 4(c, d). (1) Head and tail predictions. We also observed the imbalanced phenomenon of left-side and right-side link predictions in embedding $\mathbb{K}_{T}$. For KG2E and TransE with different temporal links, predicting the head generally yields worse MR scores than predicting the tail. However, this imbalance becomes less evident compared to $\mathbb{K}_{S}$, which suggests a more uniform structure of the temporal unit. (2) Temporal context dependency. The MR values demonstrate a consistent decrease as the time buffers increase from Past[10] to Past[60] for all cases of embedding $\mathbb{K}_{T}$. This pattern underscores the importance of including historical information to enhance link prediction accuracy in $\mathbb{K}_{T}$. Moreover, integrating more temporal links, i.e., Link[HDW], consistently bolsters MR scores for both models across all configurations, reinforcing the findings in Section IV-C.
V-B Forecasting traffic speed in limited datasets
To explore CKG-GNN’s efficacy in scenarios with potentially limited data, we apply the KGGCN models to two small datasets in Table VI, i.e., one from a typical workday $V_{work}$ and the other from a rest day $V_{rest}$. The divergence in day types is expected to manifest itself in the model’s performance due to the inherent differences in traffic patterns between workdays and rest days. The absence of weekly and daily temporal links in these single-day datasets leads to that only hourly links in $\mathbb{K}_{T}$ are utilized.
In Table VI, the CKG-enhanced models display superiority in speed prediction for both workday and rest day according to the consistently lower average MAE and MAPE for 10-120 min compared to the baseline DCRNN model. This phenomenon showcases CKG’s ability to model domain-specific context information for limited datasets, resulting in enhanced prediction accuracy. We find that a divergence in prediction ease between the workday and rest day is apparent. The $\mathbb{K}_{ST}$ model in 10-120 min demonstrates a lower MAE for the workday at 4.07±0.04 compared to 4.38±0.17 for the rest day. This discrepancy is attributed to the more structured traffic patterns typically observed during workdays, driven by routine commutes and schedules. Interestingly, when evaluating the incremental benefits offered by CKG-enhanced models, the results indicate a more pronounced improvement for the rest day than the workday. In detail, the average MAE improvement for $V_{work}$ and $V_{rest}$ is 0.16 (4.23-4.07) and 0.41 (4.79-4.38), respectively. This suggests that the integration of CKG is particularly effective in improving the less predictable and more variable traffic patterns that occur on rest days. On the rest day, the absence of regular commuting behaviors introduces complexities that context-aware knowledge graphs can offset, leading to a higher margin of improvement over the baseline.
TABLE VI: Prediction evaluation of CKG-enhanced models for two limited speed datasets. (w) refers to the workday dataset for $V_{work}$, and (r) refers to the rest day dataset for $V_{rest}$. Each experiment was run five times to obtain its mean and standard deviation.
Model 10 min 60 min 120 min Ave (10-120 min) MAE MAPE MAE MAPE MAE MAPE MAE MAPE (w)DCRNN 3.10±0.04 13.78±0.13 4.11±0.03 17.12±0.05 5.28±0.06 19.59±0.23 4.23±0.04 17.32±0.09 (w)CKG-$\mathbb{K}_{S}$ 3.08±0.01 13.51±0.05 4.10±0.01 17.00±0.03 5.13±0.04 19.41±0.14 4.18±0.02 17.15±0.05 (w)CKG-$\mathbb{K}_{T}$ 3.05±0.01 13.47±0.10 4.05±0.05 16.98±0.06 5.23±0.16 19.75±0.43 4.16±0.06 17.20±0.12 (w)CKG-$\mathbb{K}_{ST}$ 3.14±0.03 14.15±0.32 3.99±0.04 17.36±0.44 4.90±0.11 19.12±0.23 4.07±0.04 17.40±0.36 (r)DCRNN 3.27±0.06 14.01±0.51 4.51±0.04 16.33±0.05 6.46±0.14 20.33±0.29 4.79±0.05 16.97±0.12 (r)CKG-$\mathbb{K}_{S}$ 3.19±0.02 13.06±0.04 4.52±0.03 16.29±0.07 6.11±0.05 19.49±0.13 4.70±0.03 16.62±0.08 (r)CKG-$\mathbb{K}_{T}$ 3.14±0.05 12.73±0.06 4.35±0.09 15.97±0.18 5.89±0.15 19.15±0.45 4.54±0.09 16.27±0.21 (r)CKG-$\mathbb{K}_{ST}$ 3.16±0.04 13.16±0.28 4.27±0.18 15.88±0.24 5.43±0.28 17.88±0.55 4.38±0.17 16.00±0.24
Despite showing promise in limited datasets, the CKG-enhanced models also face some challenges when comparing performance across different forecast horizons. First, CKG-$\mathbb{K}_{S}$ and CKG-$\mathbb{K}_{T}$ individually outperform the baseline DCRNN model for 10-min prediction in the workday, but CKG-$\mathbb{K}_{ST}$ does not maintain this trend. We assume that when context information is focused—either spatially with $\mathbb{K}_{S}$ or temporally with $\mathbb{K}_{T}$—the models are able to effectively leverage this information for short-term predictions. When both types of context information are combined in $\mathbb{K}_{ST}$, the increased complexity requires more data to effectively train and refine its predictive capabilities in the brief 10-min forecasting interval. Second, $\mathbb{K}_{ST}$’s average performance in 10-120 min shows an improved MAE over the baseline for $V_{work}$, but its MAPE is not as favorable. The possible reasons are two-fold, i.e., relative error amplification during peak hours and lack of cyclical linkages in the single-day dataset. Since MAPE calculates the percentage error relative to the actual values, it amplifies errors when the base speed values are low in the workday peak hours. Also, MAPE’s sensitivity to errors emphasizes the challenge of accurately predicting traffic speed during congested periods, especially when lacking sufficient data samples to capture the cyclic nature of workday traffic patterns.
VI Conclusion
This study proposed an innovative context-aware knowledge graph framework designed to organize and embed spatio-temporal contexts through a relation-dependent integration strategy. Integrated with dual-view MHSA and GNN, our CKG-GNN model demonstrates state-of-the-art performance in traffic speed forecasting compared to baseline models. We find that the optimal KG configuration for the spatial unit was achieved with a buffer of [10-100] meters and [6] spatial links using ComplEx, whereas the temporal unit exhibited the best configuration with a [60] minute buffer and [Hour-Day-Week] temporal links using KG2E. These findings provide a solid foundation for employing a relation-dependent integration strategy to produce context-aware representations and capture the relationships between the traffic system and the surrounding environment. Incorporating these context-aware representations into traffic forecasting, the CKG-GNN model outperformed benchmarks, achieving an average MAE of 3.46±0.01 and a MAPE of 14.76±0.09% across speed predictions from 10 to 120 minutes. The feature-view attention heatmap highlighted the dominance of relation-dependent road features over road-only features, asserting the effectiveness of our relation-dependent integration strategy. We also observed the significance of various temporal contexts and the substantial impact of spatial and temporal links on improving forecasting accuracy. Moreover, the sequential-view attention heatmap illustrated the model’s capacity to prioritize recent time slots for predicting traffic speed. These insights affirm the utility of the proposed CKG framework in improving traffic speed forecasting, combining the strengths of knowledge graphs with neural network architectures to advance predictive performance.
While this study marks significant strides, we also encountered several challenges that may pave the way for future enhancements. First, optimize KG embedding methods. Our exploration of various KGE techniques, though comprehensive, suggests potential complexity. Building on the investigation in this study, a future enhancement could be employing AutoML techniques, such as AutoSF [36], to autonomously design scoring functions for CKGs and optimize spatio-temporal context configuration. Second, incorporate additional contexts. While our model integrates numerous spatial and temporal contexts, other influential factors that affect traffic forecasting are not included, such as road configuration. Broadening the CKG framework to include other relevant contexts could further enrich the model’s predictive capabilities. However, this expansion introduces challenges related to data integrity, processing, and the meaningful integration of new contextual features. Third, framework versatility and applicability. The CKG framework’s model-agnostic nature offers flexibility in integrating or substituting diverse time-series prediction models, which could be exploited in future research to enhance its applicability and effectiveness. Moreover, the potential of this framework extends beyond traffic forecasting. It could be adapted to other domains in intelligent transportation systems, such as informing decisions in traffic safety management or route planning.